The breaking stress of steel is 7.9×10^8 N/m^2 and density is 7.9×10^3 kg/m^3.what should be the maximum length of a stellar wire so that it may not break under it`s own weight ?

Grade:11

3 Answers

Arun
6 years ago
Dear Preeti
 
Breaking stress (W/ A) = L * rho * g/2
7.9 *10^8 = L * 7.9 * 10^3 * 10/2
2* 10^4 = L
 
 
Regards
Arun(askIITians forum expert)
Preeti
6 years ago
Breaking stress = L*rho*gL= 7.9*10^8÷7.9*10^3*9.8L=0.102*10^5L=10.2* 10^3L=10.2km So the maximum length of the wire is 10.2km
Akanksha gupta
5 years ago
BS= F/A 
BS= Mg/A....where Density(D)=M/V  
                                   V=A×L so... mass= Density×A×L 
BS= rho*A*L*g/A 
To find length, 
L= BS / D*g  
L= 7.9× 10^8/ 7.9×10^3×9.8 
L= 0.102 × 10^5 
L= 10.2km